OpenStudy (anonymous):

A friend gave me a riddle... which involves solving for x in the equation: 2^x=3^(x/2)+1. Apparently this can be solved using simple algebra, but I cannot seem to figure this out. Can anyone help?

OpenStudy (ash2326):

Is this your question? \[\huge {2^{x}=3^{\frac{x}{2}}+1}\]

OpenStudy (anonymous):

Yes. I know the answer is two, but I am trying to solve this algebraically.

OpenStudy (anonymous):

I was thinking about using the Lambert W function, but I am not familiar with that sort of thing.

OpenStudy (ash2326):

It can be written as \[\huge{e^{\ln {2^x}}=e^{\ln {3^{\frac{x}{2}}}}+1}\] Let's try by using expandion of e^x

OpenStudy (bahrom7893):

no it can't ash. If u raise both sides to e, u'll get: e^(ln2^x) = e*e^(ln(3^(x/2)))

OpenStudy (ash2326):

Bahrom I've not raised to e I've used the property \[a^{log_{a} b}=b\]

OpenStudy (bahrom7893):

but.. but u can't do that lol. how?

OpenStudy (ash2326):

That's a property of logarithms \[ \log_{2} 4= 2\] so \[2^{\log_{2} 4}=2\]

OpenStudy (ash2326):

Sorry 4

OpenStudy (bahrom7893):

wait 2^x = 3^(x/2) + 1 taking ln on both sides gives: ln(2^x) = ln(3^(x/2) + 1) u can't distribute ln in on the right.

OpenStudy (ash2326):

Nope

OpenStudy (bahrom7893):

yea so how did u get that 1 by itself on the right, that's impossible.

OpenStudy (jamesj):

Simple algebra? It's clear by simple algebra that x = 2 is a solution. What's not so obvious is that it's the only solution. To show that, you need some calculus: Consider the function f(x) = 2^x - ( 3^(x/2) + 1 ) Then f is monotonically increasing (i.e., f' > 0) and hence x = 2 can be the only solution.

OpenStudy (phi):

you can put it in the form (3+1)^y = 3^y+1 where y= x/2 expand the binomial, cancel the highest and lowest term only solution to give zero is (0 choose 1) so y=1, and x=2

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (jamesj):

@phi: that assumes that y is an integer, which it need not be.

OpenStudy (phi):

good point. Insert, for all integer y... If you want the fractional solutions, good luck finding them.